OpenStudy (anonymous):

Find the smaller solution for 3/5x^2−7=21 .

OpenStudy (anonymous):

Is this the question?: 3/(5x^2) -7 = 21

OpenStudy (anonymous):

no , 3over5xto the second power minus 7 equals 21

OpenStudy (anonymous):

so like: 3/(5x^2 - 7) = 21 ?

OpenStudy (anonymous):

yeah i guess

OpenStudy (anonymous):

are you sure? It is important.

OpenStudy (anonymous):

yeah its like that

OpenStudy (anonymous):

So first I would multiply both sides by (5x^2 - 7): 3 = 21(5x^2 - 7) then expand the left side: 3 = 105x^2 - 147 then add 147 to each side: 150 = 105x^2 then divide both sides by 105: 150/105 = x^2 Is that as far as they want you to go with it? Or did I still misunderstand the question?

OpenStudy (anonymous):

EDIT: expand the right side*

OpenStudy (anonymous):

yeah , its not like that :/

OpenStudy (anonymous):

is it like \[3/5 x^2 - 7 = 21\] ?

OpenStudy (anonymous):

yes !!!!!! :) lol

OpenStudy (anonymous):

ok so add 7 to each side: (3/5)x^2 = 21 - 7 (3/5)x^2 = 14 divide both sides by (3/5): x^2 = 14/(3/5) x^2 = 14 * (5/3) x^2 = 70/3 Hows that?

OpenStudy (anonymous):

what 70/3 what does that mean ?

OpenStudy (anonymous):

/ means division when you divide 70 by 3 you get 23.33....

OpenStudy (anonymous):

Then you find the squar root ?

OpenStudy (anonymous):

are you solving for x? if you are then you take the square root of both sides: x =about 4.83 (it can also be -4.83 but I don't know if you're expected to know that)

OpenStudy (anonymous):

yes thats the answer thank you

OpenStudy (anonymous):

no problem. do you understand how i found it?

OpenStudy (anonymous):

yea i got it

OpenStudy (anonymous):

great